OpenStudy (anonymous):

solve for t e^(t^(2))=25

OpenStudy (anonymous):

?

OpenStudy (anonymous):

e^t^2=25

OpenStudy (campbell_st):

take the base e log of both sides \[\ln(e^a) = a\] so then t^2 = ln(25) I'll let you finish it.

OpenStudy (anonymous):

did your lecture teach you ln also

OpenStudy (anonymous):

natural log

OpenStudy (anonymous):

No, section 6.1 & 6.2 are on the test this Monday and he didnt teach us any of this

OpenStudy (anonymous):

@campbell_st Do i square root both sides?

OpenStudy (anonymous):

ln works like the log,remeber for logs the base is ten \[\log a=\log_{10}a\] e a constant is a special base with a special log called the natural log so use it the same way you use logs

OpenStudy (campbell_st):

that would be the next step

OpenStudy (anonymous):

\[t=\pm \sqrt{\ln 25}\]

OpenStudy (anonymous):

hmm... thats what i did but WileyPLUS is showing that i got it incorrect

OpenStudy (anonymous):

...nevermind I had to change it up a bit thanks again